OpenStudy (anonymous):

- 1/5 - (-1) how do i answer this (the fraction is negative )

OpenStudy (anonymous):

\[-\frac{ 1 }{5 }-\left( -1 \right)=-\frac{ 1 }{5 }+1=\frac{ -1+5 }{5 }=?\]